a) The amount of money each member of a band earns playing at a graduation party varies inversely as the number of members in the band. If the band has five members, each member earns 70 dollars. Write an equation that models the relationship between the number of members in a band, n, and the amount each member earns, d.

b) Use the equation to calculate the amount each member earns if there are four members in the band.

Answer by stanbon(48558)

) The amount of money each member of a band earns playing at a graduation party varies inversely as the number of members in the band. If the band has five members, each member earns 70 dollars. Write an equation that models the relationship between the number of members in a band, n, and the amount each member earns, d.
d = k/n
Need to find "k":
70 = k/5
k = 350
----------
Equation to be used to work part b:
d = 350/n
---------------
b) Use the equation to calculate the amount each member earns if there are four members in the band.
d = 350/4
d = $87.50
=================
Cheers,
Stan H.